DOUBLE RESONANCE OVERTONE SPECTROSCOPY OF HYDRAZOIC ACID

\underline{MICHAEL R. WEDLOCK}, BERND KUHN and THOMAS R. RIZZO, Laboratoire de Chimie-Physique Moléculaire, École Polytechnique Fédérale de Lausanne, CH-1015 Lausanne, Switzerland.

We present double resonance overtone photofragment spectra of hydrazoic acid (HN_{3}) for several vibrational bands near the lowest energy spin-allowed product channel (HN_{3} → N_{2} + NH(a^{1}Δ). Double resonance overtone photofragment spectroscopy is a three-laser technique which combines IR-visible double resonance excitation of a vibrational level above the dissociation energy with LIF detection of the photofragments. We have obtained results for the fifth overtone of the NH stretch (6ν_{1}), as well as for several close-lying combination bands. With these spectra we can establish more precise limits on the dissociation threshold for the lowest spin-allowed product channel than were previously available. In addition, comparison of the underlying structure in individual rotational lines from different vibrational bands shows dramatic variations.